OpenStudy (anonymous):

What is the expression equivalent to (x^1/3)^-3

OpenStudy (btaylor):

When you have an exponent term to another exponent, you multiply the exponents. For example, \(\left(2^3\right)^2 = 2^6\). How can you use this in your problem?